A car travels 7km to east and then 4km in a direction 60° north of east.Find the magnitude and direction of the Car's displacement vector relative to positive x axis.


1
Expert's answer
2021-10-18T15:05:39-0400

Let's first find "x"- and "y"-components of the car's resultant displacement:


"d_x=7\ km\cdot cos0^{\circ}+4\ km\cdot cos60^{\circ}=9\ km,""d_y=7\ km\cdot sin0^{\circ}+4\ km\cdot sin60^{\circ}=3.46\ km."


We can find the magnitude of the car's resultant displacement from the Pythagorean theorem:


"d=\sqrt{d_x^2+d_y^2}=\sqrt{(9\ km)^2+(3.46\ km)^2}=9.64\ km."


We can find the direction of the resultant displacement from the geometry:


"\theta=tan^{-1}({\dfrac{d_y}{d_x}}),""\theta=tan^{-1}({\dfrac{3.46\ km}{9\ km}})=21^{\circ}\ N\ of\ E."